OpenStudy (anonymous):

What is the principle value of tan^−1(0) I'm working on trigonometry and I don't know where else to really look for the answer.

OpenStudy (anonymous):

\[0 \le \tan^{-1}(0) \le \frac{\pi}{2}\] http://www.efunda.com/math/trig_functions/principal.cfm

OpenStudy (anonymous):

Thanks for the answer and the link!